OpenStudy (anonymous):

Suppose you make 90% of your free throws and you attempt 3 free throws. Use binomial theorem to calculate each probability. a. You do not make any of them b. You only make 1 of them c. You only make 2 of them d. You make all of them

OpenStudy (anonymous):

C

OpenStudy (uri):

How? @sangya21 Please explain.

OpenStudy (anonymous):

It can't be C. It is four parts of one problem that I have to solve but I don't know how

OpenStudy (anonymous):

I am sorry.I thought it was an option question plus 90% of 3 = 2.7 as you are not getting 100% so you cant have all three

OpenStudy (anonymous):

That's okay @sangya21 can you help me with the problem?

OpenStudy (anonymous):

Yeah just need a little time

OpenStudy (anonymous):

Binomial formula \[P(k\space success\space for \space n \space trials) = \left(\begin{matrix}n \\ k\end{matrix}\right) *p^k*q^{n-k}\]

OpenStudy (anonymous):

Can you walk me through it?

OpenStudy (anonymous):

n = no. of trials k = success p = success probability = 0.9 q = failure = 0.1 1 n= 3; k = 0 ; p = 0.9 and q = 0.1 so P = ?

OpenStudy (anonymous):

You just said p=0.9

OpenStudy (anonymous):

its given that you find success 90% of time. so thats success probability

OpenStudy (anonymous):

I need to find the probability of making 1, 2, and 3 of the free throws.

OpenStudy (anonymous):

1. \[ P\space = \left(\begin{matrix}3 \\ 0\end{matrix}\right) (0.9)^0 (0.1)^{3-0}\] \[P\space= 3!/0!(3-0)! * 1* (0.1)^3\] \[P\space= 3!/3! * 1* (0.001)\] \[P\space= (0.001)\]

OpenStudy (anonymous):

Are you sure that's what it is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for a part. when you have 0 success

OpenStudy (anonymous):

What about b?

OpenStudy (anonymous):

in b put k = 1

OpenStudy (anonymous):

You got it?

OpenStudy (anonymous):

so is it \[P=\left(\begin{matrix}3 \\ 1\end{matrix}\right)(0.9)^{1}(0.1)^{3-1}\]

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

\[P=3!/1!(3-1)!*1*(0.1)^{3}\]

OpenStudy (anonymous):

What is next @sangya21

OpenStudy (anonymous):

0.9^1 = 0.9 so you have \[P \space = 3!/2! *0.9 *0.01\]

OpenStudy (anonymous):

P= 0.027?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

And c is \[P=\left(\begin{matrix}3 \\ 2\end{matrix}\right)(0.9)^{2}(0.1)^{3-2}\]

OpenStudy (anonymous):

\[P=3!/2!(3-2)!*1*(0.1)^{3}\]

OpenStudy (anonymous):

\[P=3!/2!*0.9*0.001\]

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

\[P = 3!/2!*1! *(0.9)^2 *0.1\] \[P = 3 *(0.9)^2 *0.1\]

OpenStudy (anonymous):

\[P = 3*0.81 *0.1\] \[P = 0.243\]

OpenStudy (anonymous):

d part \[P = (3!/3!*0!)*(0.9)^3 *(0.1)^0\] \[P = 1* (0.9)^3 *1\] because 3!/3! = 1 and (0.1)^0 = 1 \[P = 0.729\]